A financial analyst has found out that, policyholders are four times as likely to file two
claims as to file four claims. If the number of claims filed has a Poisson distribution,
what is the expected value of the number of claims filed?
1
Expert's answer
2020-06-08T19:53:13-0400

Let X be the random number of claims filed by policyholders. We know that the variance of a Poisson distribution is λ. But we have 4P(X = 4) = P(X = 2) that is 4e4/4!)=e2 /2!) ⇒4λ2/4!=1/2 ⇒ λ2 = 3⇐⇒ λ =1.732 hence the variance of X is equal to 1.732